OpenStudy (anonymous):

Explain how to find the equation of a line given two points. Do so first in words and then with an example.

OpenStudy (anonymous):

Take the difference of corresponding x and y points and then take there quotient to find the slope of the linear equation, after this substitute in an x point and a y point into your equation and solve for the unknown constant term.

OpenStudy (anonymous):

For example the line intersecting the two points $$(0,0)$$ and $$(2,10)$$ Has a slope of $$\frac{10-0}{2-0}=5$$

OpenStudy (anonymous):

So we know its of the form $$y=5x+b$$

OpenStudy (anonymous):

But $$(0,0)$$ lies on the line so we know that $$0=5*0+b=b$$

OpenStudy (anonymous):

So we now know the equation is $$y=5x$$

OpenStudy (anonymous):

There is your homework fatty copy it down

HanAkoSolo (jamierox4ev3r):

you would use the following formula\[Y^2-Y^1 \over X^2-X^1\]

HanAkoSolo (jamierox4ev3r):

lol jack so mean

OpenStudy (anonymous):

You should use subscripts not superscripts, to avoid miss communication due to ambiguous notation.

OpenStudy (anonymous):

$$\frac{y_2-y_1}{x_2-x_1}$$

HanAkoSolo (jamierox4ev3r):

lol i hate caps lock srry